OpenStudy (anonymous):

Need help to understand this ( I apparently am not understanding something as I keep getting it wrong) Find the limit. Use l'Hospital's Rule if appropriate. Use INF to represent positive infinity, NINF for negative infinity, and D for the limit does not exist. Lim t-->0 (8 e^t -8)/t^3 Step 1: verify if indetermine form: because my denominator gives me 0, I consider it an indeterminate form Step 2: I take the first derivative giving me (8 e^t )/3t^2 still gives me 0 in the denominator so I go to the second derivative Step 3: (8 e^t )/6t

OpenStudy (anonymous):

I get 0 again so I do the third derivative giving me (8 e^t )/6 and I plug in 0 in t giving me 8/6. no? but the answer is infinity

OpenStudy (anonymous):

But you got an 8 in the numeration 8/0 is not an indeterminate form

OpenStudy (anonymous):

so wait, you're telling me when I use l'hospital rule , i need a 0/0 to apply it ? or is that example just an exception ? and I thought anything over 0 would give me an error, therefore being an inderminate form

OpenStudy (anonymous):

not necessarily 0/0 there is other forms such as \[\infty^0\] \[\frac{\infty}{\infty}\] etc.. your book should show all of them... however \[\frac{8}{0}\] is not one of them

OpenStudy (anonymous):

my teacher didn't assign any books, so I work with internet , etc... I work with what i find :/ but thanks, so any number ( except 0 ) /0 is not inderminate. have a good night!